40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sat Jul 8 17:09:50 2017, gautam, Update, General, ETMY TRANS QPD anomaly ETMY_QPD.pdfETMY_QPD.png
    Reply  Sun Jul 9 08:58:07 2017, rana, Update, General, ETMY TRANS QPD anomaly 
Message ID: 13101     Entry time: Sat Jul 8 17:09:50 2017     Reply to this: 13102
Author: gautam 
Type: Update 
Category: General 
Subject: ETMY TRANS QPD anomaly 

About 2 weeks ago, I noticed some odd behaviour of the LSC TRY data stream. Its DC value seems to be drifting ~10x more than TRX. Both signals come from the transmission QPDs. At the time, we were dealing with various CDS FE issues but things have been stable on that end for the last two weeks, so I looked into this a bit more today. It seems like one particular channel is bad - Quadrant 4 of the ETMY TRANS QPD. Furthermore, there is a bump around 150Hz, and some features above 2kHz, that are only present for the ETMY channels and not the ETMX ones.

Since these spectra were taken with the PSL shutter closed and all the lab room lights off, it would suggest something is wrong in the electronics - to be investigated.

The drift in TRY can be as large as 0.3 (with 1.0 being the transmitted power in the single arm lock). This seems unusually large, indeed we trigger the arm LSC loops when TRY > 0.3. Attachment #2 shows the second trend of the TRX and TRY 16Hz EPICS channels for 1 day. In the last 12 hours or so, I had left the LSC master switch OFF, but the large drift of the DC value of TRY is clearly visible.

In the short term, we can use the high-gain THORLABS PD for TRY monitoring.

Attachment 1: ETMY_QPD.pdf  116 kB  | Hide | Hide all
ETMY_QPD.pdf
Attachment 2: ETMY_QPD.png  36 kB  Uploaded Sat Jul 8 18:15:35 2017  | Hide | Hide all
ETMY_QPD.png
ELOG V3.1.3-